OpenStudy (anonymous):

I need help with part B of this word problem A theater has 46 seats in the first row, 50 seats in the second row, 54 seats in the third row and so on. a) Can the number of seats in each row be modeled by an arithmetic or geometric sequence. b) Write the general term for the sequence a.n that gives the number of seats in row n. c) How many seats are in row 20?

OpenStudy (anonymous):

@some_someone

OpenStudy (anonymous):

46, 50, 54, ... It would be an Arithmetic Sequence.

OpenStudy (anonymous):

I know that this is an arithmetic sequence and that there are 122 seats in row 20 but I need help with B

OpenStudy (anonymous):

Do you know the common difference, d?

OpenStudy (anonymous):

Common difference is 4

OpenStudy (anonymous):

d = 54 - 50 = 4 d = 50 - 46 = 4

OpenStudy (anonymous):

Yes, good.

OpenStudy (anonymous):

Now use the formula: \[a _{n} = a _{1} = (n - 1)d\] where: \[a_{1}\] is the first term. and d is the common difference.

Directrix (directrix):

Last night, another Asker and I worked this problem with different numbers. Part B is where we snagged up. I am posting the link in the event you want to look at what we did. Again, the numbers are different but the concepts are the same. http://openstudy.com/users/directrix#/updates/5160c868e4b06dc163b9aa03

OpenStudy (anonymous):

oops \[a _{n} = a _{1} + (n - 1)d\]

OpenStudy (anonymous):

its ^ that i messed up on the "+" sorry

OpenStudy (anonymous):

\[a _{n} = 46 + (n -1)4\]

OpenStudy (anonymous):

\[a _{n} = a _{1} + (n - 1)d\] \[a _{n} = 46 + (n - 1)4\] \[a _{n} = 46 + 4n - 4 \] \[a _{n} = 4n + 42\]

OpenStudy (anonymous):

I was correct on part c right? there are 122 seats in row 20

OpenStudy (anonymous):

ok wait up :)

OpenStudy (anonymous):

\[a _{n} = a _{1} + (n - 1)d\] \[a _{20} = a _{1} + (n - 1)d\] \[a _{20} = 46 + (20 - 1)4\] \[a _{20} = 46 + (19)4\] \[a _{20} = 46 + 76\] \[a _{20} = 122\]

OpenStudy (anonymous):

yes you were :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Your Welcome :)

OpenStudy (anonymous):

Can you help me with this? \[\sum_{k= 1}^{4}(k + 2)^2\]

OpenStudy (anonymous):

@some_someone

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

I put the equation above ^^^^

OpenStudy (anonymous):

Yeah but what do you want to find? The sum?

OpenStudy (anonymous):

Sorry, Write the terms of the series and find their sum

OpenStudy (anonymous):

First, I need to ask you what you think it is?

OpenStudy (anonymous):

I dont know to be honest. I havent had a lot of practice with this.

OpenStudy (anonymous):

There is two ways to approach this. 1.) To use the properties. 2.) use some simple math

OpenStudy (anonymous):

So you know that k is the staring number, thus it says that k = 1

OpenStudy (anonymous):

and 4 is n, thus the stopping number.

OpenStudy (anonymous):

So it is the sum of \[( k + 2)^2\] from 1 to 4.

OpenStudy (anonymous):

so we just plug in any number between 1 & 4?

OpenStudy (anonymous):

So when k = 1 \[(1+2)^2\] \[(3)^2\] \[9\]

OpenStudy (anonymous):

so when k = 2 \[(k+2)^2\] \[(2+2)^2\] \[(4)^2\] \[16\]

OpenStudy (anonymous):

so when k = 3 \[(k+2)^2\] \[(3+2)^2\] \[(5)^2\] \[25\]

OpenStudy (anonymous):

So when it asks what the terms of the series are: it is k = 1, n = 4? And the sum is 9, 16, 25, 36

OpenStudy (anonymous):

so when k = 4 \[(4+2)^2\] \[(6)^2\] \[36\]

OpenStudy (anonymous):

So the sum is: 9 + 16 + 25 + 36 = 86

OpenStudy (anonymous):

oh yea right, sum, so we need to add them.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Are you good with it?

OpenStudy (anonymous):

You could also have used the properties and formulas.

OpenStudy (anonymous):

yes I am, thank you